OpenStudy (anonymous):

please help me find: the integral of sqrt(1 + 4x^2) using hyperbolic substitution

OpenStudy (anonymous):

i think the trick is to factor the 4 out of the whole thing yes?

OpenStudy (anonymous):

get \[\sqrt{4({\frac{1}{4}+x^2)}}\]

OpenStudy (anonymous):

then write \[2\sqrt{\frac{1}{4}+x^2}\]

OpenStudy (anonymous):

and the substitution you use for \[\sqrt{a^2+x^2}\] is \[x=a\sinh(u)\]

OpenStudy (anonymous):

then you do a bunch of mess about rewriting as cosh(u) and then substiting back much easier to look in a table of integrals and see that \[\int\sqrt{a^2+x^2}dx = \frac{x}{2}\sqrt{a^2+x^2}+\frac{a^2}{2}\ln(x+\sqrt{a^2+x^2})+C\]

OpenStudy (anonymous):

hahaa finally got it, thanks !

OpenStudy (anonymous):

welcome. sorry i was not much help